2) Compute the temperature rise of the ocean surface if the full solar irradiance (1380 W·m-2) is absorbed in the first 10 m of the ocean for a duration of 12 hours. Assume the first 10 m of the ocean is well mixed. The heat capacity of water is 4218 J·°C-1·Kg-1. For the density of seawater, use 1025 kg·m-3.
Answer #1

specific volume, V= 10 (m3/m2)

mass per unit area = V*density = 10*1025 = 10250 kg/m2

heat per unit area, = mcp*dT

time, t = 12 hours = 12*60*60 = 43200 seconds

heat per unit area per unit time = mcp*dT/t = 10250*4218*dT/43200

=1000.8*dT (J/ s-m2)

and given irradiation

1380 = 1000.8dT

dT = 1.378 oC